# Universal Definitions of the Roman Factorial: Introduction to Foundational Functions and the Generalization Process

### v2

@@ Acknowledgements
The author has conceived the ideas, made the calculations and has also written and approved the manuscript. This work was submitted to arXiv.org on $\pi$ day: March 14, 2024.
@@ 9.1 Number sets
A set is a mathematical concept representing a collection of distinct items, called elements or members [4] [5]. Some sets are so significant in mathematics that they have special names and notations to identify them. These important sets are often denoted using blackboard bold typeface (e.g. $\mathbb{Z}$). Common number sets include the following sets:
@@
$\mathbb{N}$ Natural numbers 1, 2, 3, 4
@@
$\mathbb{Z}$ Integers -2, -1, 0, 1
$\mathbb{Q}$ Rational numbers $\nicefrac{{1}}{{2}}$, -$\nicefrac{{5}}{{4}}$, 0.01
$\mathbb{R}$ Real numbers $\sqrt{2}$, $\pi$, $e$, $\phi$
$\mathbb{I}$ Imaginary numbers $i$, $2.5i$, $-\nicefrac{{i}}{{25}}$
$\mathbb{C}$ Complex numbers $1+i$, $\sqrt{3}-1.7i$
@@
$\mathbb{R}\setminus\mathbb{Z^{-}}\thinspace,$ (Eq. 9.1.1)
@@
Additionally, a number set without a collection of elements can be expressed in detail. For example, the set of all real numbers except 5 and 9, is:
$\mathbb{R}\setminus\{5,9\}$ (Eq. 9.1.2)
It should be noted that $\mathbb{C}$ includes all real and imaginary numbers. For instance, the number $2-3i$ is itself a complex number, as well as $5.2i$ and $4$.
9.2 $\prod$ product
The product operator $\prod$ for a product of a sequence is represented by the capital Greek letter pi, similar to how $\Sigma$ is used as the summation symbol [6]. For example, the product of the first 6 squares of natural numbers can be expressed as:
$\prod_{k=1}^{6}k^{2}=1\cdot 4\cdot 9\cdot 16\cdot 25\cdot 36$ (Eq. 9.2.1)
@@ 9.2 $\prod$ product
For the $\prod$ product to be well-defined, it’s typically required that both the upper and lower limits be integers, often natural numbers.
@@
$\prod_{k=5}^{5}k=5\thinspace,\enspace\prod_{k=3}^{3}e^{k}=e^{3}.$ (Eq. 9.2.2)
@@
$\prod_{k=3}^{2}k=\prod_{k=3}^{2}2k=\prod_{k=3}^{2}k^{3}=1\thinspace.$ (Eq. 9.2.3)
@@
$\prod_{k=n}^{n}f(k)=f(n)\thinspace,\enspace\displaystyle\prod_{k=n}^{n-a}f(k)=1\thinspace,\enspace a\in\mathbb{N}\thinspace,\enspace n\in\mathbb{Z}\thinspace,$ (Eq. 9.2.4)
@@ 9.3 Falling and rising factorials
The falling factorial (also known as the descending factorial, falling sequential product, or lower factorial [7]) is defined as the polynomial
$(x)_{n}=\prod_{k=0}^{n-1}(x-k)=x(x-1)(x-2)\vbox{\hbox{ \ldots}}(x-n+1).$ (Eq. 9.3.1)
The rising factorial (also known as the Pochhammer function, Pochhammer polynomial, ascending factorial, rising sequential product, or upper factorial [7]) is defined as
$x_{(n)}=\prod_{k=0}^{n-1}(x+k)=x(x+1)(x+2)\vbox{\hbox{ \ldots}}(x+n-1).$ (Eq. 9.3.2)
@@ 9.4 The function $\delta(n)_{f}$
$\delta(n)=\lfloor n\rfloor+\varepsilon\thinspace,\enspace 0<\varepsilon<1\thinspace,\enspace n\in\mathbb{R}\thinspace.$ (Eq. 9.4.1)
@@
$n=\lfloor n\rfloor+\{n\}\thinspace,\enspace 0\leq\{n\}<1\thinspace.$ (Eq. 9.4.2)
@@
$5.7=\lfloor 5.7\rfloor+\{5.7\}=5+0.7$
$6=\lfloor 6\rfloor+\{6\}=6+0$
$-0.2=\lfloor-0.2\rfloor+\{-0.2\}=-1+0.8$
$-3.9=\lfloor-3.9\rfloor+\{-3.9\}=-4+0.1$
It’s important to note that the floor function of a number, as defined here, is the largest integer less than or equal to that number. For negative numbers, taking the floor involves rounding down to the next smaller integer. The fractional part always lies between 0 and 1 in this context.
@@
$\lfloor x\rfloor=max\{m\in\mathbb{Z}\enspace|\enspace m\leq x\}$ (Eq. 9.4.3)
@@
Now, let’s explore the Fourier approximation of the sawtooth function. This approximation, denoted $\{n\}_{f}$ with the subscript $f$ for Fourier, is defined as:
$\{n\}_{f}=\frac{1}{2}-\frac{1}{\pi}\sum_{k=1}^{\infty}\frac{\sin(2\pi kn)}{k}\thinspace,\enspace n\in\mathbb{R}\thinspace.$ (Eq. 9.4.4)
@@
$\lfloor n\rfloor_{f}=n-\{n\}_{f}=n-\frac{1}{2}+\frac{1}{\pi}\sum_{k=1}^{\infty}\frac{\sin(2\pi kn)}{k}\thinspace,\enspace n\in\mathbb{R}\thinspace.$ (Eq. 9.4.5)
@@
$\delta(n)_{f}=\lfloor n\rfloor_{f}+\varepsilon\thinspace,\enspace n\in\mathbb{R}\thinspace,$ (Eq. 9.4.6)
@@
Given this analysis, we conclude that $\varepsilon\in(\nicefrac{{1}}{{2}},1)$, and selecting $\varepsilon=\nicefrac{{3}}{{4}}$ yields the following expression for $\delta(n)_{f}$:
$\delta(n)_{f}=n+\frac{1}{4}+\frac{1}{\pi}\sum_{k=1}^{\infty}\frac{\sin(2\pi kn)}{k}\thinspace,\enspace n\in\mathbb{R}\thinspace.$ (Eq. 9.4.7)
@@
Consequently, $\delta(n)_{f}$ can be used in place of $\delta(n)$ interchangeably. Either definition of Eq. 3.2.3 and Eq. 9.4.7 will suffice, since the F.F. $\theta(n)$ will still have the output pattern $[-1,1,1]$.
@@ 9.5 Boolean algebra
Boolean algebra [8] is a branch of algebra named after George Boole, who introduced it in his book The Mathematical Analysis of Logic (1847). It is a formal way of describing logical operations, like logic gates which are used in computing.
@@
A Boolean-valued function [9] (sometimes called a predicate or a proposition) is a function whose input is an arbitrary number set and has outputs in a Boolean domain (i.e. a generic two-element set, for example $\{0,1\}$). Those elements are interpreted as logical values, for example, $0=false$ and $1=true$ (a single bit of information).
@@
However, a few F.F. are considered Boolean-valued functions: for instance, $\xi(n)$ is a function that outputs the true/false states of the criterion $n\geq 0$. The function $\Theta(n)$, introduced in Subsection 5.2, can be thought of as an operation between two Boolean-valued functions:
@@
$\Rightarrow\enspace\Theta(n)=\xi(n)\cdot\xi(-n)$
@@ 9.6 The indeterminate form $0^{0}$
The mathematical expression $0^{0}$, known as zero to the power of zero, presents a unique case in mathematics where its definition as either 1 or left undefined depends on the context [3]. In algebra and combinatorics, it is conventionally defined as $10^{0}=$ 1. However, in mathematical analysis, this expression is often left undefined.
@@
Generally, $0^{0}$ is defined to be 1 so that many identities and properties stay true for $n=0$. One way of overcoming the anomaly is to use the F.F. $\Psi(n)$ in the base, which is never 0. Thus, an expression like $\Psi(n)^{n}$ is well defined for $n\in\mathbb{R}$.
The second limit in the first row is credited to a blackpenredpen’s video, titled finally $0^{0}$ approaches 0 after 6 years!.
@@ 9.7 The case of $0!$
The factorial of 00 is defined as $1$, denoted as $0!=1$ [10]. This definition is justified by several reasons:
@@
  • Defining $0!=1$ ensures that various combinatorial identities hold true for all valid parameter values. For example, the number of ways to select all $n$ elements from a set of $n$ is $\binom{n}{n}=\frac{n!}{n!0!}=1$, a binomial coefficient identity valid only if $0!=1$.
  • With $0!=1$, the factorial’s recurrence relation remains valid for $n=1$, simplifying recursive factorial computations by needing only the base case of zero.
@@
We chose to avoid using $10^{0}=$ 1 in this paper since it wasn’t necessary in any generalization.

### v7

@@ Acknowledgements
The author has conceived the ideas, made the calculations and has also written and approved the manuscript. The first version of this work was submitted to arXiv.org on $\pi$ day: March 14, 2024.
Software Used
This document was made in  using the online platform Overleaf, compiled in Live 2023. All figures were drawn in Mathematica. ChatGPT 3.5 has assisted in the formatting of this document.
Code Availability
The Mathematica code used to generate the figures and perform the calculations in this document is available upon request. The   document is available for downloading on arXiv.org.
For any questions, feedback, error correction or further discussion, feel free to contact me at this email: lliponis@physics.auth.gr.
@@ 9.1 Number sets
A set is a mathematical concept representing a collection of distinct items, called elements or members [setwiki] [setfun]. Some sets are so significant in mathematics that they have special names and notations to identify them. These important sets are often denoted using blackboard bold typeface (e.g. $\mathbb{Z}$). Common number sets include the following sets:
@@
$\mathbb{N}\,,\mathbb{Z}^{+}$ Natural numbers 1, 2, 3, 4 $\cdots$
@@
$\mathbb{Z}$ Integers $\cdots$ -2, -1, 0, 1 $\cdots$
$\mathbb{Q}$ Rational numbers $\nicefrac{{1}}{{2}}$, -$\nicefrac{{5}}{{4}}$, 0.01 $\cdots$
$\mathbb{R}$ Real numbers $\sqrt{2}$, $\pi$, $e$, $\phi$cdots
$\mathbb{I}$ Imaginary numbers $i$, $9.7i$, $-\nicefrac{{i}}{{25}}$cdots
$\mathbb{C}$ Complex numbers $1+i$, $\sqrt{3}-6i$cdots
@@
Note that $\mathbb{C}$ includes all real and imaginary numbers. For instance, the number $2-3i$ is itself a complex number, as well as $5.2i$ and $4$.
@@
$\mathbb{R}\setminus\mathbb{Z^{-}}\,,$
@@
Additionally, a number set without a collection of elements can be expressed in detail. For example, the set of all positive odd integers except 1, is:
$\mathbb{Z}^{+}_{odd}\setminus\{1\}\,.$
9.2 $\prod$-product
The product operator $\prod$ for a product of a sequence is represented by the capital Greek letter "pi" ($\Pi$), similar to how $\sum$ is used as the summation symbol [product]. For example, the product of the first 6 squares of natural numbers can be expressed as:
$\prod_{k=1}^{6}k^{2}=1\cdot 4\cdot 9\cdot 16\cdot 25\cdot 36\,.$
@@ 9.2 $\prod$-product
For the $\prod$-product to be well-defined, it’s typically required that both the upper and lower limits be integers, often natural numbers.
@@
$\prod_{k=5}^{5}k=5\,,\enspace\prod_{k=3}^{3}e^{k}=e^{3}.$
@@
$\prod_{k=3}^{2}k=\prod_{k=3}^{2}2k=\prod_{k=3}^{2}k^{3}=1\,.$
@@
$\prod_{k=n}^{n}f(k)=f(n)\,,\enspace\displaystyle\prod_{k=n}^{n-a}f(k)=1\,,\enspace a\in\mathbb{Z}^{+},\enspace n\in\mathbb{Z}\,,$ (Eq. 9.2.1)
@@ 9.3 Falling and rising factorials
The falling factorial (also known as the descending factorial, falling sequential product, or lower factorial [fallrisefact]) is defined as the polynomial
$(x)_{n}=\prod_{k=0}^{n-1}(x-k)=x(x-1)(x-2)\cdots(x-n+1)\,.$ (Eq. 9.3.1)
The rising factorial (also known as the Pochhammer function, Pochhammer polynomial, ascending factorial, rising sequential product, or upper factorial [fallrisefact]) is defined as
$x_{(n)}=\prod_{k=0}^{n-1}(x+k)=x(x+1)(x+2)\cdots(x+n-1)\,.$ (Eq. 9.3.2)
@@ 9.4 The function $\delta(n)_{f}$
$\delta(n)=\lfloor n\rfloor+\varepsilon\,,\quad 0<\varepsilon<1\,,\quad n\in\mathbb{R}\,.$ (Eq. 3.2.3)
@@
$n=\lfloor n\rfloor+\{n\}\,,\quad 0\leq\{n\}<1\,.$ (Eq. 9.4.1)
@@
$5.7=\lfloor 5.7\rfloor+\{5.7\}=5+0.7\,,$
$6=\lfloor 6\rfloor+\{6\}=6+0\,,$
$-0.2=\lfloor-0.2\rfloor+\{-0.2\}=-1+0.8\,,$
$-3.9=\lfloor-3.9\rfloor+\{-3.9\}=-4+0.1\,.$
It’s important to note that the floor function of a number, as defined here, is the largest integer less than or equal to that number. For negative numbers, taking the floor involves rounding down to the next smaller integer (further from 0). The fractional part always lies between 0 and 1 in this context.
@@
$\lfloor x\rfloor=max\{m\in\mathbb{Z}\enspace|\enspace m\leq x\}\,.$ (Eq. 9.4.2)
@@
Now, let’s explore the Fourier approximation of the sawtooth function [titchmarsh1986theory]. This approximation, denoted $\{n\}_{f}$ with the subscript $f$ for Fourier, is defined as:
$\{n\}_{f}=\frac{1}{2}-\frac{1}{\pi}\sum_{k=1}^{\infty}\frac{\sin(2\pi kn)}{k}\>,\quad n\in\mathbb{R}\,.$ (Eq. 9.4.3)
@@
$\lfloor n\rfloor_{f}=n-\{n\}_{f}=n-\frac{1}{2}+\frac{1}{\pi}\sum_{k=1}^{\infty}\frac{\sin(2\pi kn)}{k}\;,\enspace\>n\in\mathbb{R}\,.$ (Eq. 9.4.4)
@@
$\delta(n)_{f}=\lfloor n\rfloor_{f}+\varepsilon\,,\quad n\in\mathbb{R}\,,$ (Eq. 9.4.5)
@@
Given this analysis, we conclude that $\varepsilon\in(\nicefrac{{1}}{{2}},1)$. Selecting $\varepsilon=\nicefrac{{3}}{{4}}$ yields the following expression for $\delta(n)_{f}$:
$\delta(n)_{f}=n+\frac{1}{4}+\frac{1}{\pi}\sum_{k=1}^{\infty}\frac{\sin(2\pi kn)}{k}\>,\quad n\in\mathbb{R}\,.$ (Eq. 9.4.6)
@@
Consequently, $\delta(n)_{f}$ can be used in place of $\delta(n)$ interchangeably. Either definition of Eq. 3.2.3 and Eq. 9.4.6 will suffice, since the F.F. $\theta(n)$ will still have the output pattern $[-1,1,1]$ and $\delta(n)$ does not have any other use.
@@ 9.5 Boolean algebra
Boolean algebra [boolean_algebra] is a branch of algebra named after George Boole, who introduced it in his book The Mathematical Analysis of Logic (1847). It is a formal way of describing logical operations, like logic gates which are used in computing.
@@
A Boolean-valued function [boolean_functions] (sometimes called a predicate or a proposition) is a function whose input is an arbitrary number set and has outputs in a Boolean domain (i.e. a generic two-element set, for example $\{0,1\}$). Those elements are interpreted as logical values, for example, $0=false$ and $1=true$ (a single bit of information).
@@
However, a few F.F. are considered to be genuine Boolean-valued functions: for instance, $\xi(n)$ is a function that outputs the true/false states of the criterion $n\geq 0$. The function $\Theta(n)$, introduced in Subsection 5.2, can be thought of as an operation between two Boolean-valued functions:
@@
$\Rightarrow\enspace\Theta(n)=\xi(n)\cdot\xi(-n)\,.$
@@ 9.6 The indeterminate form $0^{0}$
The mathematical expression $0^{0}$, known as zero to the power of zero, presents a unique case in mathematics where its definition as either 1 or left undefined depends on the context [0tothe0]. In algebra and combinatorics, it is conventionally defined as $10^{0}=$ 1. However, in mathematical analysis, this expression is often left undefined.
@@
Generally, $0^{0}$ is defined to be 1 so that many identities and properties stay true for $n=0$. One way of overcoming the anomaly is to use the foundational function $\Psi(n)$ in the base, which is never 0. Thus, an expression like $\Psi(n)^{n}$ is well defined for $n\in\mathbb{R}$. We chose to avoid using $10^{0}=$ 1 in this paper since it wasn’t necessary in any F.F. or generalization.
In the examples above, the second limit in the first row is credited to a blackpenredpen’s video, titled finally $0^{0}$ approaches 0 after 6 years!.
@@ 9.7 The case of $0!$
The factorial of 00 is defined as $1$, denoted as $0!=1$ [0_factorial]. This definition is justified by several reasons:
@@
[…]
  • With $0!=1$, the factorial’s recurrence relation remains valid for $n=1$, simplifying recursive factorials by needing only the base case of zero.
@@
In this study, we do not strive to prove why/if relationships like $0!=1$ are true, instead we try to find intuitive reasons for their validity in this context.